A 1 liter solution contains 0.495 M hydrofluoric acid and 0.371 M sodium fluoride.

Addition of 0.408 moles of hydrobromic acid will:
(Assume that the volume does not change upon the addition of hydrobromic acid.)

A. Raise the pH slightly

B.Lower the pH slightly

C. Raise the pH by several units

D. Lower the pH by several units

E. Not change the pH

F. Exceed the buffer capacity
Answer #1

HF (aq.) + H2O (l) = H3O+ (aq.) + F- (aq.)

NaF (aq.) ---------------> Na+ (aq.) + F- (aq.)

According to Henderson's and Hessalbatch's equaiton of buffers to calculate pH,

pH = pKa + Log[salt]/[acid]

Additon of strong acid i.e hydrobromic acid, produces H+ ions. These ion react with anions of salt to form unionised weak acid as follows,

H+ (aq.) + F- (aq.) = HF (aq.)

Because of this reaction the concentration of acid will increase and salt will decrease.

From the formul provided concentration of acid in inversely proportional to pH. Hence increase in concentration of acid will decrease the pH.

So, (B) Lower the pH slightly.